# Trainwreck: A damaging adversarial attack on image classifiers

### v1

@@ 4.1 Stealth
In line with Section 3, Trainwreck treats stealth as a satisficing objective, setting three strict stealth requirements:
@@
Figure 2: Examples of two pairs of clean and perturbed images ($\epsilon\leq\frac{8}{255}$) with varying, but still acceptable levels of stealth.
The total number of images in the poisoned dataset (Equation 3) and the number of images for each class (Equation 4) must match the respective counts in the original training dataset. Data count is a simple, easily accessible, and often used statistic—altering it is a dead giveaway that something is amiss with the data. Equation 5 ensures the stealth of individual adversarial perturbations, for each restricting its strength, given by its $\ell^{\infty}$-norm, to a small $\epsilon$ [9]. Inspired by existing literature [10, 27, 40], we explicitly set $\epsilon=\frac{8}{255}$. This value corresponds to up to 8 pixel intensity level difference per channel in the 8-bit color encoding. Figure 2 shows examples of clean images with their perturbed counterparts with $\epsilon\leq\frac{8}{255}$. Despite light, homogenous backgrounds like in the ship example on the right tending to highlight the perturbations, the overall demonstrated stealth level is acceptable.
@@ 4.2 Estimating class distribution divergence
Trainwreck has full access to the data, but the attacker cannot be assumed to know which classes are similar to each other. Therefore, Trainwreck must estimate class similarity automatically, performing a heuristic estimate of distribution divergence between all pairs of classes. As a preliminary step, Trainwreck extracts an auxiliary feature representation $\bar{\mathbf{X}}_{tr}$ of dimension $n_{tr}\times n_{feat}$ from $\mathcal{D}_{tr}$. The objective is to obtain generic discriminative features of “just good enough” quality. Therefore, it is sufficient to use a publicly available model pre-trained on a natural image dataset. As features, the default choice for Trainwreck is the output of the ViT-L-16 vision transformer [19] with pre-trained ImageNet SWAG weights available in the torchvision PyTorch package.
@@ 4.3 Class-pair universal perturbations
With the stealth rules set and the class distances estimated, there is one final preliminary step: to train a surrogate model $f_{s}$ that Trainwreck uses to craft adversarial perturbations. The surrogate model is trained on $\mathcal{D}_{tr}$ and evaluated on $\mathcal{D}_{ts}$. The better the surrogate model’s test accuracy, the better it understands the data and therefore the better adversarial perturbations it is able to produce. Trainwreck is a black-box attack, so the choice of the surrogate model’s architecture does not matter: as long as it has reasonable test accuracy, Trainwreck should work well. This claim is verified by the experimental evaluation in Sections 5 and 6.
@@
              if $f_{s}(\mathbf{X}_{tr}^{(i)}+\tilde{\mathbf{\Delta}}^{<c_{a}>})=c_{a}$ then
@@
                  $\tilde{\mathbf{\Delta}}^{<c_{a}>}\leftarrow\tilde{\mathbf{\Delta}}^{<c_{a}>}+\tilde{\mathbf{\Delta}}^{(i)}$
                  $\tilde{\mathbf{\Delta}}^{<c_{a}>}\leftarrow\textsc{Clip}(\tilde{\mathbf{\Delta}}^{<c_{a}>},-\epsilon,\epsilon)$
@@
Algorithm 1 Method for calculating class-based universal perturbations (CPUPs).
@@ 4.4 Trainwreck algorithm
              $\mathbf{X}_{tr}^{(i)}\leftarrow\mathbf{X}_{tr}^{(i)}+\tilde{\mathbf{\Delta}}^{<c_{a}>}$
@@
Trainwreck first completes the preliminary steps: extracts the features, computes the class divergence matrix $\mathbf{D}$ (Section 4.2), and trains the surrogate model $f_{s}$ used to craft the CPUPs (Section 4.3). Trainwreck attacks all classes in the dataset. For each attacked class $c_{a}$, the method first determines which images of $c_{a}$ are attacked and which stay clean. Trainwreck employs a simple strategy: it poisons the same ratio of images for each class, equal to the global poison rate $\pi$. For each class $c_{a}$, it uniformly randomly selects $\lfloor\pi n^{<c_{a}>}\rfloor$ images to be poisoned, leaving the remaining $\lceil(1-\pi)n^{<c_{a}>}\rceil$ images untouched. The index set of the poisoned images is denoted $\tilde{I}$. Next, the closest class $c_{c}$, $c_{c}\neq c_{a}$, with the minimum divergence from $c_{a}$ according to $\mathbf{D}$ is selected. Then, the CPUP for $c_{a}$ is computed. Note that the computation of CPUP uses all the training data of class $c_{a}$, regardless of $\pi$. Finally, all images in $\tilde{I}$ get updated with the CPUP.
The output is the poisoned training dataset $\mathcal{D}_{tr}$ that for each attacked class $c_{a}$ shifts $\lfloor\pi n^{<c>}\rfloor$ images towards its closest class $c_{c}$ using the CPUP, leaving the rest of the data and all the class labels intact. The method satisfies all three stealth criteria outlined in Section 4.1: the number of images in the dataset and per class remains the same, and the perturbation strength restriction is strictly adhered to.
@@ 5 Experimental setup
Since DAAs have to the best of our knowledge not been systematically explored yet, there are no existing benchmarks. This section therefore proposes a challenging experimental setup to evaluate Trainwreck’s claimed benefits. In particular, it aims to answer the following questions:
  • (EQ1) Is Trainwreck an effective DAA?
  • (EQ2) Can Trainwreck be replaced by intuitive baselines?
  • (EQ3) Is Trainwreck’s strength customizable?
  • (EQ4) Is Trainwreck indeed black-box and transferable?
@@
Method EfficientNetV2 ResNeXt FT-ViT EfficientNetV2 ResNeXt FT-ViT
@@
Table 1: Test top-1 accuracy results of the strongest versions of the evaluated attacks.
5.1 Baselines
The experimental evaluation uses three DAA baselines. Two of them are swapper attacks that swap up to $\lfloor\frac{\pi n}{2}\rfloor$ training images between different classes, which ensures the poison rate $\pi$ is not exceeded. The first swapper attack is RandomSwap, which simply performs uniformly randomly selected swaps. The second is JSDSwap, a more “informed” swapper that uses the $\mathbf{D}$ matrix of Jensen-Shannon divergence (JSD) values (see Section 4.2). JSDSwap goes down the list of classes ranked in ascending order by their minimal JSD from other classes, from each swapping $\lfloor\frac{\pi n^{<c>}}{2}\rfloor$ uniformly randomly selected images with images from the closest class according to $\mathbf{D}$.
Figure 3: Comparison of 2x4 image grids displaying CIFAR-10 data of class dog poisoned by Trainwreck and a swapper attack (here: RandomSwap), both with $\pi=0.25$.
The final baseline, AdvReplace, is a perturbation-based attack that simply replaces a fraction of the training data proportional to $\pi$ with adversarial images resulting from running an untargeted PGD attack on the surrogate model. Comparing Trainwreck with JSDSwap and AdvReplace essentially completes an ablation study: JSDSwap tests the importance of class divergence knowledge, AdvReplace tests adversarial perturbations without class knowledge.
5.2 Datasets and parameters
The attack methods have been evaluated on CIFAR-10 and CIFAR-100 [20], both well-established image classification benchmarks. The CIFAR datasets are simple: it is quite easy for modern models to learn to classify them them high accuracy. This is precisely what makes them a challenging benchmark for DAAs, as it is hard to confuse the models with small perturbations or with swapping a few images.
The perturbation strength parameter $\epsilon$ is set to $\frac{8}{255}$ to satisfy the stealth requirements set in Section 4.1. The poison rate $\pi$ does not need to change between datasets, but it does vary between swapper and perturbation attacks. Swapper attacks are visually conspicuous, explicitly breaking the perturbation stealth rule (see Equation 5), as their perturbation is the pixel difference between the swapped image and the original. As shown in Figure 3, the same $\pi$ results in completely different levels of stealth between perturbation and swapper attacks. With $\pi=0.25$, Trainwreck or AdvReplace would likely fly under the radar, while a swapper attack is easily noticeable. To balance this effect out, we iterate over $\pi\in[0.05,0.25]$ in swapper attack experiments, treating the situation depicted in Figure 3 as an upper bound w.r.t. stealth. In the perturbation attack experiments, we iterate over $\pi\in[0.25,1]$: $\pi=1$ is permissible due to the full adherence to the stealth requirements, while $\pi<0.25$ would result in minimal to no effect on the target model due to the perturbations being small. For the other parameters, we use the empirical values as described in Section 1: $n_{iter}^{cpup}=1$, $n_{iter}^{pgd}=10$.
5.3 Evaluation protocol
To evaluate Trainwreck and the baselines, we perform attacks on two datasets and three target model architectures. Each evaluated attack method first poisons the training dataset $\mathcal{D}_{tr}$, creating a poisoned dataset $\tilde{\mathcal{D}_{tr}}$. Then, the target model is trained on $\tilde{\mathcal{D}_{tr}}$. For each evaluated method, we report the attacked model’s test top-1 accuracy on the original test dataset $\mathcal{D}_{ts}$ (note that none of the attack methods tamper with $\mathcal{D}_{ts}$). To eliminate early-epoch randomness, the highest accuracy value in the final 10 epochs is reported. The damage to test accuracy is a measure of the evaluated DAA’s success, in line with Section 3.
The three evaluated target model architectures are EfficientNetV2 [33], ResNeXt-101 [37], and the ViT-L-16 vision transformer [19]. As the surrogate model, Trainwreck and AdvReplace use ResNet-50 [12]. All EfficientNetV2 and ResNeXt-101 models were trained from scratch, all ViT-L-16 models had been initialized to pre-trained ImageNet weights and finetuned to the CIFAR datasets: only the new fully connected output layer is trained, the rest of the parameters are frozen. Therefore, the ViT models are further referred to as FT-ViT. All models have been trained using the PyTorch framework, each model is the respective standard implementation from the torchvision package. All models have been trained for 30 epochs: this is not enough to reach full performance, but more than enough for the effects of the evaluated attacks to converge.
This choice of models allows evaluating a number of scenarios. Firstly, the surrogate model’s architecture is less advanced than the targets’: in a real setting, the attacker will likely have a weaker model than the target due to lower domain knowledge. This allows for testing realistic transferability. To put the black-box claim to the test, one of the target model architectures, ResNeXt, is more similar to the surrogate model than the rest. If the results on ResNeXt are noticeably better than on the other models, then it is doubtful that the evaluated attack is black-box. Finally, the difficulty of attacking FT-ViT should not be understated: it entails manipulating a large model with $\sim$300M parameters, pretrained on a rich, general image dataset, while only being able to poison a single output layer.
6 Experimental results
@@
Table 1 reports the test top-1 accuracy for the strongest versions of the evaluated attacks, i.e., with $\pi=1$ for the perturbation attacks and $\pi=0.25$ for the swapper attacks. For the sake of completeness, the test top-1 accuracy of the surrogate ResNet-50 model is 0.8557 on CIFAR-10 and 0.5923 on CIFAR-100, both lower than the respective target models. Figure 4 plots test top-1 accuracy for varying $\pi$.
The answer to EQ1 is that Trainwreck indeed is a highly effective DAA. On both model architectures trained from scratch, EfficientNetV2 and ResNeXt, it is far ahead of the competition in reducing the target model’s test accuracy. Within 30 epochs, the target models attacked by Trainwreck reach a performance barely above random choice and stay there. On the difficult FT-ViT model, the accuracy drop caused by Trainwreck is several percent (1.95% on CIFAR-10 and 7.75% on CIFAR-100). This accuracy drop is smaller, but a noticeable adversarial effect is still present. Trainwreck can therefore outright break models trained from scratch on the poisoned dataset, and still leave a dent in a pre-trained “monolith” ViT model where it could only influence the final output layer parameters with stealthy perturbations.
To answer EQ2, overall, the intuitive baselines are not an adequate substitute for Trainwreck, although they can be a better choice in certain scenarios. In particular, the FT-ViT model seems to respond more strongly to swapper attacks than to Trainwreck. If the attacker can afford the stealth loss, opting for a swapper attack may be a solid choice. Looking at the results closer, a likely explanation is that the stealthy perturbations of Trainwreck simply do not have the “gradient power” to sway the mostly-frozen FT-ViT model during finetuning. After all, the strongest attack on FT-Vit, JSDSwap, uses the same class divergence knowledge as Trainwreck. The deciding factor therefore are JSDSwap’s more drastic data edits. On all other models, Trainwreck is the clearly superior choice.
Figure 4 gives insight into the customizability of Trainwreck’s strength (EQ3). For both datasets, the relationship between the test accuracy and $\pi$ can be decomposed roughly into three regions. For $\pi<0.5$, the attacks have a limited effect, $\pi\in[0.5,0.8]$ results in a noticeable, moderate drop in performance, and for $\pi>0.8$, the performance starts degrading rapidly. Selecting $\pi$, which is a free choice for the attacker, therefore determines the attack strength, making Trainwreck a customizable attack.
To answer EQ4: the black-box claim is verified by the absence of a performance spike on ResNeXt-101. The surrogate model with a similar architecture does not have a noticeable advantage, i.e., the attacker does not have to match the target model’s architecture. Transferability is verified by the single poisoning of each dataset resulting in a degraded performance across target models. The experiments do indicate some variance in attack strength depending on the model type—however, as explained in previous text, it is more difficult to attack FT-ViT with perturbation attacks than the other models by the very nature of FT-ViT. Trainwreck therefore is a black-box and transferable attack.

### v2

@@ 4.1 Stealth
Figure 2: Examples of two pairs of clean and perturbed images ($\epsilon\leq\frac{8}{255}$).
To meet the stealth DAA requirement defined in Section 3, Trainwreck treats stealth as a satisficing objective, composed of three sub-objectives:
@@
The total number of images in the poisoned dataset (Equation 3) and the number of images for each class (Equation 4) must match the respective counts in the original training dataset. Data count is a simple, easily accessible, and often used statistic—altering it is a dead giveaway that something is amiss with the data. Equation 5 ensures the stealth of individual adversarial perturbations, for each restricting its strength, given by its $\ell^{\infty}$-norm, to a small $\epsilon$ [9]. Inspired by existing literature [10, 29, 40], we explicitly set $\epsilon=\frac{8}{255}$. Figure 2 shows examples of clean images with their perturbed counterparts with $\epsilon\leq\frac{8}{255}$. Despite light, homogenous backgrounds like in the ship example on the right tending to highlight the perturbations, the overall demonstrated stealth level is acceptable.
@@ 4.2 Estimating class distribution divergence
In general, the attacker cannot be assumed to know which classes are similar. Therefore, Trainwreck must estimate class similarity automatically, performing a heuristic estimate of distribution divergence between all pairs of classes. As a preliminary step, Trainwreck extracts an auxiliary feature representation $\bar{\mathbf{X}}_{tr}$ of dimension $n_{tr}\times n_{feat}$ from $\mathcal{D}_{tr}$. The objective here is to obtain “just good enough” features, so it is sufficient to use a publicly available model pre-trained on a natural image dataset. Our default choice are output features of the ViT-L-16 vision transformer [21] with pre-trained ImageNet SWAG weights.
@@ 4.3 Class-pair universal perturbations
There is one final preliminary step: to train a surrogate model $f_{s}$ that Trainwreck uses to craft adversarial perturbations. The surrogate model is trained on $\mathcal{D}_{tr}$ and evaluated on $\mathcal{D}_{ts}$. The better the surrogate model’s test accuracy, the better it understands the data and therefore the better adversarial perturbations it is able to produce. Trainwreck is a black-box attack, so the choice of the surrogate model’s architecture does not matter: as long as it has reasonable test accuracy, Trainwreck should work well. This claim is verified by the experimental evaluation in Sections 5 and 6.
@@
              if $\textsc{Clip}(f_{s}(\mathbf{X}_{tr}^{(i)}+\tilde{\mathbf{\Delta}}^{<c_{a}>}),-\epsilon,\epsilon)=c_{a}$ then
@@
                  $\tilde{\mathbf{\Delta}}^{<c_{a}>}\leftarrow\textsc{Clip}(\tilde{\mathbf{\Delta}}^{<c_{a}>}+\tilde{\mathbf{\Delta}}^{(i)},-\epsilon,\epsilon)$
@@
Algorithm 1 Calculating class-based universal perturbations (CPUPs).
@@ 4.4 The full attack
              $\mathbf{X}_{tr}^{(i)}\leftarrow\textsc{Clip}(\mathbf{X}_{tr}^{(i)}+\tilde{\mathbf{\Delta}}^{<c_{a}>},-\epsilon,\epsilon)$
@@
In the preliminary steps, Trainwreck extracts the features, computes the class divergence matrix $\mathbf{D}$ (Section 4.2), and trains the surrogate model $f_{s}$ used to craft the CPUPs (Section 4.3). Trainwreck attacks all classes in the dataset. For each attacked class $c_{a}$, the method first determines which images of $c_{a}$ are attacked and which stay clean. Trainwreck employs a simple strategy: it poisons the same ratio of images for each class, equal to the global poison rate $\pi$. For each class $c_{a}$, it uniformly randomly selects $\lfloor\pi n^{<c_{a}>}\rfloor$ images to be poisoned, leaving the remaining $\lceil(1-\pi)n^{<c_{a}>}\rceil$ images untouched. The index set of the poisoned images is denoted $\tilde{I}$. Next, the closest class $c_{c}$, $c_{c}\neq c_{a}$, with the minimum divergence from $c_{a}$ according to $\mathbf{D}$ is selected. Then, the CPUP for $c_{a}$ is computed. Note that the computation of CPUP uses all the training data of class $c_{a}$, regardless of $\pi$. Finally, all images in $\tilde{I}$ get updated with the CPUP.
The output is the poisoned training dataset $\mathcal{D}_{tr}$ that for each attacked class $c_{a}$ shifts $\lfloor\pi n^{<c>}\rfloor$ images towards its closest class $c_{c}$, leaving the rest of the data and all the class labels intact. The method satisfies all three stealth criteria outlined in Section 4.1: the number of images in the dataset and per class remains the same, and the perturbation strength restriction is always met.
@@ 5 Experimental setup
Bearing in mind the DAA requirements set in Section 3, we answer the following questions to experimentally evaluate Trainwreck’s performance:
  • (EQ1) Is Trainwreck a potent DAA?
  • (EQ2) Is Trainwreck a customizable DAA?
  • (EQ3) Is Trainwreck black-box and transferable?
  • (EQ4) Ablation study: Can Trainwreck be replaced by a subset of its components?
Note that stealth, the third and final DAA requirement, is not evaluated experimentally, as Trainwreck assures stealth by strictly adhering to the stealth objectives outlined in Section 4.1.
5.1 Baselines
The experimental evaluation pits Trainwreck against seven baselines in total: three state-of-the-art data poisoning methods, three custom DAA baselines, and a random guess model (Random) as a reference frame. The three existing data poisoning approaches are learning to confuse (L2Confuse) by Feng et al. [6], adversarial poisoning (AdvPoison) by Fowl et al. [8], and unlearnable examples (UnlearnEx) by Huang et al. [15]. All three recent data poisoning baselines use adversarial perturbation poison and operate under similar stealth requirements as Trainwreck, providing for a fair and challenging comparison
Since the existing data poisoning attacks cannot be compared to Trainwreck beyond EQ1, to provide more detailed results in DAA context and flesh out the answers to the remaining EQs, we introduce three custom baselines. Two of them are swapper attacks that swap up to $\lfloor\frac{\pi n}{2}\rfloor$ training images between different classes, which ensures the poison rate $\pi$ is not exceeded. The first swapper attack is RandomSwap, which simply performs uniformly randomly selected swaps. The second is JSDSwap, a more “informed” swapper that uses the $\mathbf{D}$ matrix of Jensen-Shannon divergence (JSD) values (see Section 4.2). JSDSwap goes down the list of classes ranked in ascending order by their minimal JSD from other classes, from each swapping $\lfloor\frac{\pi n^{<c>}}{2}\rfloor$ uniformly randomly selected images with images from the closest class according to $\mathbf{D}$.
The final baseline, AdvReplace, is a perturbation-based attack that simply replaces a fraction of the training data proportional to $\pi$ with adversarial images resulting from running an untargeted PGD attack on the surrogate model. Essentially, AdvReplace can be thought of as a generic, simple adversarial perturbation based poisoning attack. Comparing Trainwreck with JSDSwap and AdvReplace completes the ablation study, testing the importance of class divergence knowledge and adversarial perturbations without class knowledge, respectively.
5.2 Datasets and parameters
The attack methods have been evaluated on CIFAR-10 and CIFAR-100 [22], both established image classification benchmarks. The CIFAR datasets are simple: it is quite easy for modern models to learn to classify them them high accuracy. Precisely this makes them a challenging benchmark for DAAs, as it is hard to confuse the models with small perturbations or with swapping a few images.
[…]
5.3 Evaluation protocol
We have run Trainwreck and custom baseline attacks on three target model architectures. For each evaluated method, we report the attacked model’s test top-1 accuracy on the original test dataset $\mathcal{D}_{ts}$ (note that none of the attack methods tamper with $\mathcal{D}_{ts}$). To eliminate early-epoch randomness, the highest accuracy value in the final 10 epochs is reported. The damage to test accuracy is a measure of the evaluated DAA’s success, in line with Section 3.
The three target model architectures evaluated in our experiments are EfficientNetV2 (EN-V2) [35], ResNeXt-101 (RN-101) [38], and the ViT-L-16 vision transformer (FT-ViT) [21]. As the surrogate model, Trainwreck and AdvReplace use ResNet-50 [12]. All EN-V2 and RN-101 models were trained from scratch, all transformer models had been initialized to pre-trained ImageNet weights and finetuned: only the new fully connected output layer is trained, the rest of the parameters are frozen. All models have been trained for 30 epochs: not enough for full performance, but sufficient for the attack effects to converge.
This choice of models allows evaluating a number of scenarios. Firstly, the surrogate model’s architecture is less advanced than the targets’: in a real setting, the attacker likely has a weaker model than the target due to lower domain knowledge. This tests realistic transferability. To test the black-box claim, one of the target architectures, RN-101, is more similar to the surrogate model than the rest. If RN-101 results are noticeably better, then it is doubtful the attack is black-box. Finally, the difficulty of attacking FT-ViT should not be understated: it entails manipulating a large model with $\sim$300M parameters, pretrained on a rich image dataset, while poisoning a single output layer.
6 Experimental results
Table 1 reports the strongest poisoning results achieved by Trainwreck and the existing state-of-the-art poisoning attacks. Table 2 reports the test top-1 accuracy for the strongest versions of Trainwreck and the custom baselines. For the sake of completeness, the test top-1 accuracy of the surrogate ResNet-50 model is 0.8557 on CIFAR-10 and 0.5923 on CIFAR-100, both lower than the respective target models. Figure 3 plots test top-1 accuracy for varying $\pi$.
Attack Surrogate arch. Target arch. CIFAR-10 CIFAR-100
Random N/A N/A 0.1 0.01
L2Confuse [6] U-Net ConvNet 0.2877 N/A
AdvPoison [8] ResNet-18 ResNet-18 0.0625 N/A
UnlearnEx [15] ResNet-18 ResNet-50 0.1345 0.0380
Trainwreck ($\pi=1)$ ResNet-50 ResNeXt-101 0.1211 0.0135
Table 1: Test accuracy of models damaged by train data poisoning. For each attack, the lowest achieved accuracy and the corresponding architectures are reported.
@@ 6 Experimental results
Method EN-V2 RN-101 FT-ViT EN-V2 RN-101 FT-ViT
@@
Table 2: Test top-1 accuracy results of Trainwreck and the custom baselines.
The answer to EQ1 is that Trainwreck indeed is a potent attack. Within 30 epochs, the target models trained from scratch on Trainwreck-attacked data permanently degrade to a performance barely above random choice. As Table 1 shows, this potency is comparable or even better than the data poisoning state of the art. On the difficult FT-ViT model, the accuracy drop caused by Trainwreck is several percent (1.95% on CIFAR-10 and 7.75% on CIFAR-100). This accuracy drop is smaller, but a noticeable adversarial effect is still present, even though the attack could only influence the final output layer parameters with stealthy perturbations.
Figure 3 convincingly answers EQ2: Trainwreck is a customizable DAA. For both datasets, the relationship between the test accuracy and $\pi$ can be decomposed roughly into three regions. For $\pi<0.5$, the attacks have a limited effect, $\pi\in[0.5,0.8]$ results in a moderate drop in performance, and for $\pi>0.8$, the performance starts degrading rapidly. Since the attacker can choose $\pi$ freely and thus influence the attack strength, Trainwreck is indeed customizable. Note that none of the data poisoning baselines are customizable: L2Confuse [6] and UnlearnEx [15] require full poisoning ($\pi=1$), the AdvPoison paper reports their partial poisoning merely “does not significantly improve results over training on only the clean data and often degrades results below what one would achieve using only the clean data” [8]. Customizability is therefore, to the best of our knowledge, unique to Trainwreck.
@@
To answer EQ3: the black-box claim is verified by the absence of a performance spike on ResNeXt-101. The surrogate model with a similar architecture does not have a noticeable advantage, i.e., the attacker does not have to match the target model’s architecture. Transferability is verified by the single poisoning of each dataset resulting in a degraded performance across target models. Some variance in attack strength is present due to FT-ViTs being difficult to attack by nature in this setting. Overall, Trainwreck is black-box and transferable.
To answer EQ4, overall, Trainwreck’s components, JSDSwap and AdvReplace, are not an adequate substitute for Trainwreck, although they can be a better choice in certain scenarios. In particular, the FT-ViT model seems to respond more strongly to swapper attacks than to Trainwreck. Looking at the results closer, a likely explanation is that the stealthy perturbations of Trainwreck simply do not have the “gradient power” to sway the mostly-frozen FT-ViT model during finetuning. The deciding factor therefore are JSDSwap’s more drastic data edits. On all other models, Trainwreck is clearly the superior choice.